Project

General

Profile

Actions

2017-06-30-LH2

30th June 2017; 11:00 BST

Phone details

Present: JB, JT, AN, KL, AK, MT, PW

Notes

Actions

  • Action AK follow through de-bugging of FC datalogger.
    • Stands. Main issue to investigate is why signal that indicates a quench is high. AK will rerun the test. Issue is recording of data in the event of a quench.
  • Action CMcW to install NI on MICEnet machine.
    • Stands. We were not sure whether this had been done. So, action stands.
  • Action MT/JB to agree and implement testing of quench valves.
    • Stands. JB has slightly modified the quench valve to introduce chamfers to avoid metal-to-metal contact on return. Need to devise method of testing. To be done next week.
  • Action MT to label keys for H2 system.
    • Done.
  • Action PW to modify purge sequence limits in agreement with JB.
    • Stands.
  • Action MT/JB/AK to agree monitoring GUI.
    • Stands. Snapshot of GUI has been sent from AK to JB. JB says needs to be simplified. MT sent suggestion. AK, MT, and JB need to meet to discuss.
  • Action AK to verify read/write permission control via PLC IP address control.
    • Done. Instructions in document circulated by MC are clear. This means that AK can implement after conversation with PW. Need to clarify how HMI is connected to PLC.
  • Action AK to organize meeting to discuss alarm handler and agree archiver PVs.
    • Stands. Do at meeting to discuss GUI.
  • Action CW/JB prevent rotation of H2 connectors
    • Done.

Monitoring and archival of data - PF/AK

Current technical

  • FC pressure rating
    • quench/leak scenarios.
    • route to PV nominated engineer approval
  • T piece for bypass line.
  • Refurbishment of pipework from turret to gas panel and plenum
    • Y piece - external fabrication. JT
    • 150mm line JT
    • H2 supply line 'boot' - with manufacturer.
    • Quench line support - ?? to design?
  • Refurbishment of pipework from plenum to roof and vent
    • Swap 2nd sections and cut new holes in shed at higher level to clear obstructions - PL to implement..
    • move hydrogen insulating volume backing pump to east 500mm done CW.

Timetable

  • New turret install Wed 28th
    attachment:17-06-22PlanCW.xlsx
  • Next meeting
  • AoB

Back to LH2:

Back

Files:

Updated by Long, Kenneth almost 6 years ago · 3 revisions